### v2

@@ Abstract
In this work, we study the task of scheduling jobs on a single machine with sequence dependent family setup times under the goal of minimizing the makespan, that is, the completion time of the last job in the schedule. This notoriously NP-hard problem is highly relevant in practical productions and requires heuristics that provide good solutions quickly in order to deal with large instances. In this paper, we present a heuristic based on the approach of parameterized local search. That is, we aim to replace a given solution by a better solution having distance at most $k$ in a pre-defined distance measure. This is done multiple times in a hill-climbing manner, until a locally optimal solution is reached. We analyze the trade-off between the allowed distance $k$ and the algorithm’s running time for four natural distance measures. Example of allowed operations for our considered distance measures are: swapping $k$ pairs of jobs in the sequence, or rearranging $k$ consecutive jobs. For two distance measures, we show that finding an improvement for given $k$ can be done in $f(k)\cdot n^{\mathcal{O}(1)}$ time, while such a running time for the other two distance measures is unlikely. We provide a preliminary experimental evaluation of our local search approaches.
Keywords:
makespan minimization hill-climbing algorithms parameterized local search.
@@ 1 Introduction
Finding orderings in which products are manufactured on a machine is among the most important problems in combinatorial optimization and is highly relevant in practical production scenarios. It has applications in many real-world scenarios appearing for instance in the solar cell industry [3] or the cider industry [33]. Since the mid-1960s, scheduling problems are extensively studied for a wide range of shop floor models, target functions, and setup information [1].
@@
In this work, we study the task of scheduling jobs on a single machine with sequence dependent setup times [17, 36] under the goal of minimizing the makespan, that is, the completion time of the last job in the schedule. In this problem, one is given a set of jobs, each with an individual processing time, a deadline, and some identifier for a product type. Additionally, one is given a setup matrix specifying the additional time it takes to change the machine setup from one product type to another. The goal is to find a sequence of all jobs such that every job meets its deadline and the completion time of the last job is minimal. Our work is motivated by the real-world production scheduling of the company CUNA.
@@
The considered scheduling problem is $\mathrm{NP}$-hard as it generalizes the famous Traveling Salesperson Problem (TSP). Due to this intractability, instances with a large number of jobs are usually solved using heuristics. A common approach to tackle scheduling problems are genetic algorithms [30, 26]. Since the applications of genetic algorithms usually have a large number of parameters that play an important role in their performance [5], it can be very difficult to reconstruct studied algorithms [32] and make them produce equally good results for a new use case. Moreover, genetic algorithms usually do not provide a (local) optimality guarantee for the returned solution. A further important class of heuristics are local search algorithms, which usually have a simple design and provide a local optimality guaranty on the returned solution: In a hill climbing manner, one aims to improve a solution by performing one small change on a current solution, until no further improvement can be found. Variable Neighborhood Search (VNS) strategies that aim to find an improvement by combining swapping two jobs and inserting a job at another position in the schedule [9, 31] are considered a state-of-the-art approach for single-machine scheduling [1].
@@
Our approach is based on another version of local search called parameterized local search. While in other approaches a solution was improved by one operation [9, 31], the user may set a search radius $k$ and extend the search space for possible improvements with up to $k$ operations. More precisely, one aims to find an improving solution within a radius of size $k$ around the current solution for some distance function. This technique might prevent getting stuck in bad local optima. With this work, we aim to outline the trade-off between the running time of the heuristic and the size of the search radius $k$. Parameterized local search lately received much attention in the algorithmics community both from a theoretical [2, 8, 13, 18, 25, 34] and practical [12, 11, 16, 19, 24] point of view. In case of sequencing problems, parameterized local search has been studied for TSP [27] and for finding topological orderings of Bayesian networks [16].
@@
Parameterized local search combines the paradigms of parameterized algorithmics [6] with local search heuristics. Parameterized algorithmics for single machine scheduling has been studied extensively: The parameterized complexity of minimizing the weighted number of tardy jobs [23] has been studied, also in combination with the weighted completion time [21]. Fellows and McCartin [10] studied single-machine scheduling with precedence constraints. For the szenario where jobs can be rejected with some rejection cost, de Weerdt et al. [7] provided results on the parameterized algorithmics and studied approximation schemes. A relatively new setting is equitable singe-machine scheduling with multiple agents guaranteeing each agent a minimal level of service [20]. For a structured overview, especially on open challenges in parameterized algorithmics for scheduling problems, we refer to the survey by Minch and van Bevern [28]. While there are these many works on parameterized algorithmics for single-machine scheduling for different scenarios than the one studied in this work, this is—to the best of our knowledge—the first work studying parameterized local search for a scheduling problems.
@@
In the realm of parameterized local search, the choice of the distance and the choice of the radius $k$ in which we aim to find an improvement is crucial. We now discuss the four distances investigated in this work. When considering the insert distance, one aims to improve the solution by removing up to $k$ jobs from the ordering and inserting them at new positions. When considering the swap distance, one aims to improve the solution by sequentially swapping up to $k$ pairs of jobs in a given schedule. Note that parameterized local search for these two distances essentially lifts established local search strategies via single swaps and single inserts [9, 31] to the parameterized approach and is thus a natural next step in the design of hill-climbing strategies. When considering the window distance, one checks whether there exists a window consisting of $k$ consecutive jobs in a given schedule, such that the solution can be improved by rearranging the jobs inside the window. When considering the multi-window distance, one can rearrange the jobs in arbitrarily many disjoint windows, of up to $k$ consecutive jobs each, simultaneously to obtain an improved solution. Figure 1 shows one example of each studied distance. Note that the multi-window distance can be seen as an extension of the window distance in the following sense: The set of possible schedules with multi-window distance $k$ is a superset of the schedules with window distance $k$ from a given schedule. The multi-window distance is closely related to a multi-inversions distance used for finding Bayesian network structures via parameterized local search [16]. Note that for all these distances, the problem of searching for a better schedule with distances at most $k$ becomes $\mathrm{NP}$-hard since for sufficiently large $k$, one aims to find a globally optimal solution.
@@
Figure 1: Examples of the considered distance measures: the swap distance of $\pi$ and $\pi_{1}$ is 3, the insert distance of $\pi$ and $\pi_{2}$ is 1, the window distance of $\pi$ and $\pi_{3}$ is 4, and the multi-window distance of $\pi$ and $\pi_{4}$ is 3.
@@
We show for the window distance and the multi-window distance, that one can find in time $k^{\min{(k,t)}}\cdot|I|^{\mathcal{O}(1)}$ an improving schedule if one exists. Herein, $|I|$ denotes the total input size and $t$ denotes the number of distinct product types in the instance. Note that in this running time, there is no exponential dependence on the total input size. Since the number of different product types manufactured on one machine is usually bounded by a small constant, the factor $k^{\min{(k,t)}}$ provides a good trade-off between the running time and the size of the search radius $k$. We complement this result by showing that for the insert and the swap distance, there is no algorithm with running time $f(k)\cdot|I|^{\mathcal{O}(1)}$ for any computable function $f$, unless $\mathrm{W}[1]=\mathrm{FPT}$. While the contributions of this paper are mainly theoretical results for the parameterized local search problems, we also provide preliminary experiments to asses whether parameterized local search is in principle a viable approach in practice. More precisely, we evaluate the solution quality of our hill-climbing approaches focused on 12 test instances. Motivated by the practical relevance of genetic algorithms and by the good results of VNS strategies, we consider two genetic algorithms and a state-of-the-art hill-climbing algorithm PILS1 [31] as a base line for the comparison of our preliminary experiments.
@@ 2 Preliminaries
Given some integers $a\in\mathds{N}$ and $b\in\mathds{N}$, we let $[a,b]:=\{n\in\mathds{N}\mid a\leq n\leq b\}$. We study a problem of sequencing jobs on a single machine with sequence-dependent family setup times. Formally, a job corresponds to a triple $j:=(t_{j},d_{j},\tau_{j})$. Herein, the value $t_{j}\in\mathds{N}_{0}$ denotes the processing time of $j$, the value $d_{j}\in\mathds{N}_{0}$ denotes the deadline of $j$, and $\tau_{j}$ is an identifier of a type of $j$. Let $\mathcal{T}$ denote the set containing all possible types. A setup mapping is a mapping $\ell:\mathcal{T}\times\mathcal{T}\rightarrow\mathds{N}_{0}$. Given two types $\tau$ and $\tau^{\prime}$, the value $\ell(\tau,\tau^{\prime})$ is the time needed to set up the machine, if a job of type $\tau^{\prime}$ is scheduled right after a job of type $\tau$. We assume that $\ell(\tau,\tau)=0$ for every $\tau\in\mathcal{T}$ and that $\ell$ satisfies the triangle inequality. That is, $\ell(\tau_{1},\tau_{3})\leq\ell(\tau_{1},\tau_{2})+\ell(\tau_{2},\tau_{3})$ for every triple $(\tau_{1},\tau_{2},\tau_{3})$ of types.
@@
In the systematic Graham notation [15], MM is the problem $1\mid\text{ST}_{\text{sd,f}}\mid C_{\max}$: Scheduling on one machine, when we have sequence-dependent family setup times and aiming to minimize the makespan $C_{\max}$.
@@
A distance measure $\delta:\pi\times\pi^{\prime}\mapsto x\in\mathds{N}$ maps a pair of schedules to a non-negative integer. The distance measures considered in this work satisfy $\delta(\pi,\pi)=0$ for every $\pi$ and are symmetric, that is, $\delta(\pi,\pi^{\prime})=\delta(\pi^{\prime},\pi)$ for every $\pi$ and $\pi^{\prime}$.
@@
$\delta$ LS MM
@@
Task: Find a schedule $\pi^{\prime}$ of $J$ with $\delta(\pi,\pi^{\prime})\leq k$ such that $\pi^{\prime}$ is better than $\pi$, that is, $\pi^{\prime}$ is feasible and $C_{\max}^{\pi^{\prime}}<C_{\max}^{\pi}$.
@@ 2. $b)$
    there is some index $i\in[n-k+1,n]$ such that $\{\pi(t)\mid t\in[i+1,n]\}=\{\pi^{\prime}(t)\mid t\in[i+1,n]\}$, and the subsequences $\pi[1,i]$ and $\pi^{\prime}[1,i]$ decompose into $k$-intervals.
@@
We next consider Statement $a)$. Let $\pi$ be a feasible schedule with window distance at most $k$ from $\pi$ such that $C^{\pi^{\prime}}_{\max}<C^{\pi}_{\max}$. Furthermore, let $\xi$ be the number of type-inversions in the rearranged window $\pi^{\prime}[a,b]$, such that $\xi$ is minimal among all such $\pi^{\prime}$. Note that in the proof of $b)$, rearranging one single rearranged window $\pi^{\prime}[I_{i}]$ was sufficient to obtain a contradiction to the minimality of $\xi$ in the case of $\xi>0$. By repeating the same arguments for the single rearranged window $\pi^{\prime}[a,b]$, we conclude $\xi=0$. Thus, Statement $a)$ holds. $\square$
We would like to emphasize that the triangle inequality—which is needed for Proposition 1—is a natural probability of setup functions in practice: it appears to be unrealistic that setting up to an intermediate type $\tau^{\prime}$ accelerates setting up a machine from a type $\tau_{1}$ to another type $\tau_{2}$. Even if that was the case, one would declare this intermediate setup as the standard setup from $\tau_{1}$ to $\tau_{2}$ and consider the updated setup function instead.
@@
We next consider the running time of the algorithm. The number of table entries is the product of $|\mathcal{T}|$ and the number of possible prefix vectors $\vec{p}$. Since the values of the $\tau$th entry of such prefix vector range from 00 to $q_{\tau}$, there are at most $\prod_{\tau\in\mathcal{T}}(q_{\tau}+1)$ such prefix vectors. Since each entry can be computed in $|I|^{\mathcal{O}(1)}$ time, the algorithm has the stated running time. $\square$
@@
The correctness of the algorithm follows by the fact that we consider every combination of $(i,\tau^{\prime},\tau^{\prime\prime})$, and that it is sufficient to find the best EDDS in a rearranged window due to Proposition 2 $a)$. For the running time, note that each instance of Internal MM can be solved in $k^{\min(k,t)}\cdot n^{\mathcal{O}(1)}$ time. Since we solve at most $n\cdot t^{2}$ such instances, the algorithm has the claimed running time. $\square$
@@
We next analyze the running time of the algorithm. The dynamic programming table $T$ has $n\cdot t$ entries; recall that $t$ is the number of distinct types. To compute a single entry, we iterate over each combination of $j$$\tau^{\prime}$, and $\tau^{\prime\prime}$. Consequently, one table entry can be computed with $k\cdot t^{2}$ evaluations of the subroutine. Since Internal MM can be solved in $k^{\min{(k,t)}}n^{\mathcal{O}(1)}$ time, the algorithm has the claimed running time. $\square$

### v3

@@ Abstract
In this work, we study the task of scheduling jobs on a single machine with sequence dependent family setup times under the goal of minimizing the makespan.This notoriously NP-hard problem is highly relevant in real-world industrial productions and requires heuristics that provide good solutions quickly in order to deal with large instances. In this paper, we present a heuristic based on the approach of parameterized local search. That is, we aim to replace a given solution by a better solution having distance at most $k$ in a pre-defined distance measure. This is done multiple times in a hill-climbing manner, until a locally optimal solution is reached. We analyze the trade-off between the allowed distance $k$ and the algorithm’s running time for four natural distance measures. Examples of allowed operations for our considered distance measures are: swapping $k$ pairs of jobs in the sequence, or rearranging $k$ consecutive jobs. For two distance measures, we show that finding an improvement for given $k$ can be done in $f(k)\cdot n^{\mathcal{O}(1)}$ time, while such a running time for the other two distance measures is unlikely. We provide a preliminary experimental evaluation of our local search approaches.
@@ 1 Introduction
Finding orderings in which products are manufactured on a machine is among the most important problems in combinatorial optimization and is highly relevant in real-world industrial production. It has applications in many real-world scenarios appearing for instance in the solar cell industry [3] or the cider industry [33]. Since the mid-1960s, scheduling problems are extensively studied for a wide range of shop floor models, target functions, and setup information [1].
@@
In this work, we study the task of scheduling jobs on a single machine with sequence dependent family-setup times [17, 36] under the goal of minimizing the makespan, that is, the completion time of the last job in the schedule. In this problem, one is given a set of jobs, each with an individual processing time, a deadline, and some identifier for a product type. Additionally, one is given a setup matrix specifying the additional time it takes to change the machine setup from one product type to another. The goal is to find a sequence of all jobs such that every job meets its deadline and the completion time of the last job is minimal. Our work is motivated by the real-world production scheduling of the company CUNA. CUNA manufactures products that come in eight different material/color combinations on a single injection molding machine and changing the setup from one material/color combination to another takes a combination specific amount of time.
@@
The considered scheduling problem is $\mathrm{NP}$-hard as it obviously generalizes the famous Traveling Salesperson Problem (TSP). Due to this intractability, instances with a large number of jobs are usually solved using heuristics. A common approach to tackle scheduling problems are genetic algorithms [30, 26]. Since the applications of genetic algorithms usually have a large number of parameters that play an important role in their performance [5], it can be very difficult to reconstruct studied algorithms [32] and make them produce equally good results for a new use case. Moreover, genetic algorithms usually do not provide a (local) optimality guarantee for the returned solution. Another important class of heuristics are local search algorithms, which usually have a simple design and provide a local optimality guarantee on the returned solution: In a hill climbing manner, one aims to improve a solution by performing one small change on a current solution, until no further improvement can be found. Variable Neighborhood Search (VNS) strategies that aim to find an improvement by combining swapping two jobs and inserting a job at another position in the schedule [9, 31] are considered a state-of-the-art approach for single-machine scheduling [1].
@@
Our approach is based on another version of local search called parameterized local search. While in other approaches a solution is improved by one operation [9, 31], the user may set a search radius $k$ and extend the search space for possible improvements with up to $k$ operations. More precisely, one aims to find an improving solution within a radius of size $k$ around the current solution for some distance function. This technique might prevent getting stuck in bad local optima. With this work, we aim to outline the trade-off between the running time of the heuristic and the size of the search radius $k$. Parameterized local search lately received much attention in the algorithmics community both from a theoretical [2, 8, 13, 18, 25, 34] and practical [12, 11, 16, 19, 24] point of view. In case of sequencing problems, parameterized local search has been studied for TSP [27] and for finding topological orderings of Bayesian networks [16].
@@
Parameterized local search combines the paradigms of parameterized algorithmics [6] with local search heuristics. Parameterized algorithmics for single machine scheduling has been studied extensively: The parameterized complexity of minimizing the weighted number of tardy jobs has been studied [23], also in combination with the weighted completion time [21]. Fellows and McCartin [10] studied single-machine scheduling with precedence constraints. For the scenario where jobs can be rejected with some rejection cost, de Weerdt et al. [7] provided results on the parameterized algorithmics and studied approximation schemes. A relatively new setting is equitable singe-machine scheduling with multiple agents guaranteeing each agent a minimal level of service [20]. For a structured overview, especially on open challenges in parameterized algorithmics for scheduling problems, we refer to the survey by Minch and van Bevern [28]. While there are many works on parameterized algorithmics for single-machine scheduling for different scenarios than the one studied in this work, this is—to the best of our knowledge—the first work studying parameterized local search for a scheduling problem.
@@
In the realm of parameterized local search, the choice of the distance function on the solution space and the choice of the search radius $k$ in which we aim to find an improvement is crucial. We now discuss the four distances investigated in this work. When considering the insert distance, one aims to improve the solution by removing up to $k$ jobs from the ordering and inserting them at new positions. When considering the swap distance, one aims to improve the solution by sequentially swapping up to $k$ pairs of jobs in a given schedule. Note that parameterized local search for these two distances essentially lifts established local search strategies via single swaps and single inserts [9, 31] to the parameterized approach and is thus a natural next step in the design of hill-climbing strategies. When considering the window distance, one checks whether there exists a window consisting of $k$ consecutive jobs in a given schedule, such that the solution can be improved by rearranging the jobs inside the window. When considering the multi-window distance, one can rearrange the jobs in arbitrarily many disjoint windows, of up to $k$ consecutive jobs each, simultaneously to obtain an improved solution. Figure 1 shows one example of each studied distance. Note that the multi-window distance can be seen as an extension of the window distance in the following sense: The set of possible schedules with multi-window distance $k$ is a superset of the schedules with window distance $k$ from a given schedule. The multi-window distance is closely related to a multi-inversions distance used for finding Bayesian network structures via parameterized local search [16]. Note that for all these distances, the problem of searching for a better schedule with distances at most $k$ becomes $\mathrm{NP}$-hard, since for sufficiently large $k$, one aims to find a globally optimal solution.
@@
Figure 1: Examples of the considered distance measures: the insert distance of $\pi$ and $\pi_{1}$ is 1, the swap distance of $\pi$ and $\pi_{2}$ is 3, the window distance of $\pi$ and $\pi_{3}$ is 4, and the multi-window distance of $\pi$ and $\pi_{4}$ is 3.
@@
We show for the window distance and the multi-window distance that one can find in time $k^{\min{(k,t)}}\cdot|I|^{\mathcal{O}(1)}$ an improving schedule if one exists. Herein, $|I|$ denotes the total input size and $t$ denotes the number of distinct product types in the instance. Note that in this running time, there is no exponential dependence on the total input size. Since the number of different product types manufactured on one machine is usually bounded by a small constant, the factor $k^{\min{(k,t)}}$ provides a good trade-off between the running time and the size of the search radius $k$. We complement this result by showing that for the insert and the swap distance, there is no algorithm with running time $f(k)\cdot|I|^{\mathcal{O}(1)}$ for any computable function $f$, unless $\mathrm{W}[1]=\mathrm{FPT}$. While the contributions of this paper are mainly theoretical results for the parameterized local search problems, we also provide preliminary experiments to asses whether parameterized local search is in principle a viable approach in practice. More precisely, we evaluate the solution quality of our hill-climbing approaches on 12 test instances. Motivated by the practical relevance of genetic algorithms and by the good results of VNS strategies, we consider two genetic algorithms and a state-of-the-art hill-climbing algorithm PILS1 [31] as a base line for the comparison of our preliminary experiments.
@@ 2 Preliminaries
Given some integers $a\in\mathds{N}_{0}$ and $b\in\mathds{N}_{0}$, we let $[a,b]:=\{n\in\mathds{N}_{0}\mid a\leq n\leq b\}$. We study a problem of sequencing jobs on a single machine with sequence-dependent family setup times. Formally, a job corresponds to a triple $j:=(t_{j},d_{j},\tau_{j})$. Herein, the value $t_{j}\in\mathds{N}_{0}$ denotes the processing time of $j$, the value $d_{j}\in\mathds{N}_{0}$ denotes the deadline of $j$, and $\tau_{j}$ is an identifier of a type of $j$. Let $\mathcal{T}$ denote the set containing all possible types. A setup mapping is a mapping $\ell:\mathcal{T}\times\mathcal{T}\rightarrow\mathds{N}_{0}$. Given two types $\tau$ and $\tau^{\prime}$, the value $\ell(\tau,\tau^{\prime})$ is the time needed to set up the machine, if a job of type $\tau^{\prime}$ is scheduled right after a job of type $\tau$. We assume that $\ell(\tau,\tau)=0$ for every $\tau\in\mathcal{T}$ and that $\ell$ satisfies the triangle inequality. That is, $\ell(\tau_{1},\tau_{3})\leq\ell(\tau_{1},\tau_{2})+\ell(\tau_{2},\tau_{3})$ for every triple $(\tau_{1},\tau_{2},\tau_{3})$ of types.
@@
In the systematic Graham notation [15], MM is the problem $1\mid\text{ST}_{\text{sd,f}},\overline{d}_{j}\mid C_{\max}$: Scheduling on one machine, when we have sequence-dependent family setup times, strict deadlines, and aiming to minimize the makespan $C_{\max}$.
@@
A distance measure $\delta:\pi\times\pi^{\prime}\mapsto x\in\mathds{N}_{0}$ maps a pair of schedules to a non-negative integer. The distance measures considered in this work satisfy $\delta(\pi,\pi)=0$ for every $\pi$ and are symmetric, that is, $\delta(\pi,\pi^{\prime})=\delta(\pi^{\prime},\pi)$ for every $\pi$ and $\pi^{\prime}$.
@@
$\delta$ Local Search Makespan Minimization ($\delta$ LS MM)
@@
Task: Find a schedule $\pi^{\prime}$ of $J$ with $\delta(\pi,\pi^{\prime})\leq k$ such that $\pi^{\prime}$ is better than $\pi$, that is, $\pi^{\prime}$ is feasible and $C_{\max}^{\pi^{\prime}}<C_{\max}^{\pi}$. Or report that no better schedule exists.
@@ 2. $b)$
    there is some index $i\in[n-k,n]$ such that $\{\pi(t)\mid t\in[i+1,n]\}=\{\pi^{\prime}(t)\mid t\in[i+1,n]\}$, and the subsequences $\pi[1,i]$ and $\pi^{\prime}[1,i]$ decompose into $k$-intervals.
@@
We next consider Statement $a)$. Let $\pi$ be a feasible schedule with window distance at most $k$ from $\pi$ such that $C^{\pi^{\prime}}_{\max}<C^{\pi}_{\max}$. Furthermore, let $\xi$ be the number of type-inversions in the rearranged window $\pi^{\prime}[a,b]$, such that $\xi$ is minimal among all such $\pi^{\prime}$. Note that in the proof of $b)$, modifying one single rearranged window $\pi^{\prime}[I_{i}]$ was sufficient to obtain a contradiction to the minimality of $\xi$ in the case of $\xi>0$. By repeating the same arguments for the single rearranged window $\pi^{\prime}[a,b]$, we conclude $\xi=0$. Thus, Statement $a)$ holds. ∎
We would like to emphasize that the triangle inequality—which is needed for Proposition 1—is a natural property of setup functions in practice: it appears to be unrealistic that setting up to an intermediate type $\tau^{\prime}$ accelerates setting up a machine from a type $\tau_{1}$ to another type $\tau_{2}$. Even if that was the case, one would declare this intermediate setup as the standard setup from $\tau_{1}$ to $\tau_{2}$ and consider the updated setup function instead.
@@
We next consider the running time of the algorithm. The number of table entries is the product of $|\mathcal{T}|$ and the number of possible prefix vectors $\vec{p}$. Since the values of the $\tau$th entry of such prefix vector range from 00 to $q_{\tau}$, there are at most $\prod_{\tau\in\mathcal{T}}(q_{\tau}+1)$ such prefix vectors. Since each entry can be computed in $|I|^{\mathcal{O}(1)}$ time, the algorithm has the stated running time. ∎
@@
The correctness of the algorithm follows by the fact that we consider every combination of $(i,\tau^{\prime},\tau^{\prime\prime})$, and that it is sufficient to find the best EDDS in a rearranged window due to Proposition 2 $a)$. For the running time, note that each instance of Internal MM can be solved in $k^{\min(k,t)}\cdot n^{\mathcal{O}(1)}$ time. Since we solve at most $n\cdot t^{2}$ such instances, the algorithm has the claimed running time. ∎
@@
We next analyze the running time of the algorithm. The dynamic programming table $T$ has $n\cdot t$ entries; recall that $t$ is the number of distinct types. To compute a single entry, we iterate over each combination of $j$$\tau^{\prime}$, and $\tau^{\prime\prime}$. Consequently, one table entry can be computed with $k\cdot t^{2}$ evaluations of the subroutine. Since Internal MM can be solved in $k^{\min{(k,t)}}n^{\mathcal{O}(1)}$ time, the algorithm has the claimed running time. ∎